[Brahms-l] Shift report 20050523 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon May 23 2005 - 08:03:19 EDT
Shifter leader was rk
Shifters were 


1. Beam from previous shift dumped earlier than scheduled due to low yellow polarization (-28.2%). New store from 0300 (polarization more than 40%, now more than 50%) still on.
2. Runs 14864-14867, 14873-14881 at full field B, 2.3FS and 90MRS taken.
14871,14872 - pedestals.
3. None observed.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon May 23 08:03:31 2005

This archive was generated by hypermail 2.1.8 : Mon May 23 2005 - 08:03:44 EDT